# Timelike conjugate points in Lorentzian length spaces

### v1

@@ 1 Introduction
Our article is structured as follows: In Section 2, we first recall some basics on Lorentzian pre-length spaces used throughout (Subsection 2.1). We then study geodesics and their convergence properties in Subsection 2.2. Due to the lack of arclength parametrization in the generality that we work with, we need to consider equivalence classes of geodesics up to suitable reparametrizations. To this end, in Appendix A, we give a treatment of the Fréchet distance on such equivalence classes of non-stopping curves on a metric spaces that is suitable for our purposes. Our main result in this context is the closure of the set of (equivalence classes of) causal geodesics $\operatorname{Geo}(X)$ in the set of non-stopping curves with respect to the Fréchet distance (see Proposition 2.26).
@@ 2.1 Preliminaries on Lorentzian pre-length spaces
future-directed timelike (resp. causal) curve (past-directed curves are defined analogously) in a Lorentzian pre-length space $X$ is a nowhere constant map $\gamma\colon I=[a,b]\to X$ (i.e., $\gamma$ is not constant on any non-trivial subinterval of $I$) that is Lipschitz (w.r.t. $d$) and such that for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$ we have $\gamma(t_{1})\ll\gamma(t_{2})$ (resp.) $\gamma(t_{1})\leq\gamma(t_{2})$ (causal curves on unbounded intervals can be defined analogously). Its $\tau$-length is defined as
@@
A reparametrization of a future-directed causal curve $\gamma\colon[a,b]\to X$ is a future-directed causal curve $\lambda\colon[c,d]\to X$ such that $\gamma=\lambda\circ\varphi$, where $\varphi\colon[a,b]\to[c,d]$ is continuous, surjective, and strictly monotonically increasing. A future-directed causal curve $\gamma\colon[a,b]\to X$ is called a maximizer if $L_{\tau}(\gamma)=\tau(\gamma(a),\gamma(b))$.
In particular, the reparametrized curve $\lambda\circ\varphi$ above must itself be Lipschitz. The $\tau$-length of a causal curve is invariant under reparametrizations ([12, Lem. 2.18]). We note that causal curves were defined in [12] as non-constant maps rather than nowhere constant. We choose the current convention for two main reasons, one being that the restriction of a causal curve to any non-trivial subinterval of its domain should again be causal, and the other that causal curves in the standard smooth spacetime setting are indeed nowhere constant. The Fréchet distance as detailed in Appendix A then applies to causal curves.
@@
localizable Lorentzian pre-length space is a Lorentzian pre-length space equipped with well-behaved local time separation functions, see [12, Def. 3.16] for the precise definition. It is called regularly localizable if the maximizers w.r.t. the local time separations (which exist by definition of localizability) are required to be timelike if they connect timelike related points. Regularly localizable Lorentzian pre-length spaces are regular, cf. [12, Thm. 3.18].
@@ 2.2 Convergence of geodesics
A future-directed causal curve $\gamma\colon I\to X$ is a geodesic if for each $t\in I$ there exists a neighborhood $[a,b]$ of $t$ (i.e., $a<t<b$, allowing for equality at the endpoints of $I$) such that $\gamma|_{[a,b]}$ is a maximizer, i.e., $\tau(\gamma(a),\gamma(b))=L_{\tau}(\gamma|_{[a,b]})$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a localizable Lorentzian pre-length space. Then it is finitely $\tau$-measurable.
@@
This follows from the definition of localizability, cf. [12, Def. 3.16] for the latter. ∎
@@
An important class of examples of Lorentzian (pre-) length spaces is furnished by smooth manifolds $M$ equipped with a continuous and causally plain metric $g$ (cf. [12, Prop. 5.8]). To compare the notion of geodesics in Lorentzian manifolds with that of Definition 2.8, let us now extend the validity of [12, Prop. 2.32] from smooth Lorentzian metrics to a large class of continuous ones.
@@
To prove this we employ a sequence of smooth Lorentzian metrics $g_{k}$ with wider lightcones than $g$, approximating $g$ as constructed in [14, Lem. A.1]. Then for the corresponding time separation functions $\tau_{k}$ of $g_{k}$ we have that $\tau$ is the locally uniform and non-increasing limit of the $\tau_{k}$ (see [14, Lem. A.2]). In particular, $L_{\tau_{k}}(\gamma)\geq L_{\tau}(\gamma)$ for all $k$. Let $\varepsilon>0$ and choose a partition $P=\{a=t_{0}<\dots<t_{N}=b\}$ of $[a,b]$ such that
@@
It follows from Proposition 2.13 that (in strongly causal spaces) geodesics in the sense of Definition 2.8 will coincide with geodesics in the sense of solutions of the geodesic equation as long as the regularity of the metric is sufficient to allow for a satisfactory solution theory of these ODEs and, additionally, as long as such solutions are indeed local maximizers of the $g$-length-functional. This certainly is the case for $g\in C^{1,1}$ (cf. [15] or [13]). However, the Hartmann–Wintner-type spacetime $M\coloneqq\mathbb{R}\times(-1,1)\times\mathbb{R}^{n}$ with metric $g\coloneqq-dt^{2}+dx^{2}+(1-|x|^{\lambda})g_{n}$, where $\lambda\in(1,2)$ and $g_{n}$ is the Euclidean metric on $\mathbb{R}^{n}$, is an example of a $C^{1,\lambda-1}$-Lorentzian metric where solutions of the geodesic equations fail to even locally maximize the $g$-length, cf. [19, Ex. 2.4].
@@
Note that, unlike [10, Definitions 2.18 & 2.19], we do not require that $(X,d,\ll,\leq,\tau)$ be causally path-connected in our definition of $\leq_{U}$ and, hence, of weakly causally closed. However, we have the following slightly sharpened version of [10, Prop. 2.21]:
@@
    $(X,d,\ll,\leq,\tau)$ is locally causally closed; [^4] Recall from [12, Definition 3.4] that if $(X,d,\ll,\leq,\tau)$ is a Lorentzian pre-length space, then a subset $U\subseteq X$ is causally closed if, for all sequences $(p_{n})_{n}$, $(q_{n})_{n}$ in $U$ such that $p_{n}\leq q_{n}$ for all $n$ with $p_{n}\to p\in\bar{U}$, $q_{n}\to q\in\bar{U}$, we have $p\leq q$. A Lorentzian pre-length space is locally causally closed if every point has a causally closed neighborhood.
@@
Finally, as pointed out in [10, pp. 8], the limit curve theorem [12, Thm. 3.7] holds with the assumption of local weak causal closedness.
@@
From the proof of [12, Thm. 3.7], we deduce that there exists a subsequence $\left(\gamma_{n_{k}}\right)_{k}$ of curves that converge uniformly to a Lipschitz continuous limit curve $\gamma$. Therefore, without loss of generality, we may assume that the sequence of curves $\left(\gamma_{n}\right)_{n}$ converges uniformly to $\gamma$. Since $X$ is locally weakly causally closed, for every $t\in[a,b]$ there exists an open, weakly causally closed neighborhood $U_{t}$ of $\gamma(t)$. Let $a\leq t_{1}<t_{2}\leq b$ be such that $\gamma(t_{1}),\gamma(t_{2})\in U_{t}$. Then, by uniform convergence, there exists an $n_{0}\in\mathbb{N}$ such that, for all $n\geq n_{0}$, we have $\gamma_{n}\!\left(\left[t_{1},t_{2}\right]\right)\subseteq U_{t}$. Since $\gamma_{n}$ are causal, this implies that $\gamma_{n}(t_{1})\leq_{U_{t}}\gamma_{n}(t_{2})$ for all $n\geq n_{0}$. The fact that $U_{t}$ is weakly causally closed then implies that $\gamma(t_{1})\leq_{U_{t}}\gamma(t_{2})$ and, therefore, $\gamma(t_{1})\leq\gamma(t_{2})$. The rest of the proof of part (i) proceeds as in the proof of [12, Lemma 3.6].
@@
Let $(M,g)$ be a strongly causal $C^{1,1}$-spacetime, then it is locally maximizing: By Remark 2.11, the two available notions of timelike (null, causal) geodesics agree. Now let $p\in M$ and $U$ a convex neighborhood containing $p$. Then (cf. [13, 15]) for any two causally related points $q_{1}\leq q_{2}$, $q_{1},q_{2}\in U$, there is a unique causal geodesic connecting them which is entirely contained in $U$. Hence any causal geodesic in $U$ is already maximizing.
@@
The Hartman-Wintner-type spacetimes from Remark 2.11 are not locally maximizing.
@@
The Lorentzian constant curvature $K$ spaces, which are (the universal coverings of) de-Sitter space, Minkowski space, and anti-de Sitter space of the corresponding radius (cf. Definition 2.3), are locally maximizing: Indeed in the case of $K\geq 0$ this is trivial since timelike geodesics are maximizing throughout, and in the case $K<0$ one can consider convex neighborhoods of a given point $x$ which avoid the conjugate locus $\{\tau(x,.)=\frac{\pi}{\sqrt{-K}}\}$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space. We denote by $\operatorname{Geo}(X)$ the set of equivalence classes of future-directed causal geodesic segments $\gamma\colon[0,1]\to X$, where two such segments are called equivalent if and only if there exists a continuous, strictly increasing and surjective reparametrization $\varphi\colon[0,1]\to[0,1]$ between them. [^6] See Definition 2.2. We denote by $\operatorname{TGeo}(X)\subseteq\operatorname{Geo}(X)$ the subset of equivalence classes of future-directed timelike geodesic segments. For convenience, we write $\gamma\in\operatorname{Geo}(X)$ for both the equivalence class and some representative. We consider the following metrics on $\operatorname{Geo}(X)$: $d_{F}$ is the restriction of the Fréchet distance (see Appendix A), and
@@
Let $(X,d,\ll,\leq,\tau)$ be a localizable, locally weakly causally closed, strongly causal, finitely $\tau$-measurable and locally maximizing Lorentzian pre-length space. Then $\operatorname{Geo}(X)$ is a closed subset of the set of nowhere constant curves from $[0,1]$ to $X$ w.r.t. the Fréchet distance $d_{F}$. Moreover, the metric topologies of $d_{F}$ and $d_{\Gamma}$ coincide on $\operatorname{Geo}(X)$.
@@
Let $(\gamma_{n})\subseteq\operatorname{Geo}(X)$ converge to the equivalence class of a nowhere constant path $\gamma\colon[0,1]\to X$. By definition of $d_{F}$, there exists a sequence of representatives (again called $\gamma_{n}$) that converge uniformly to $\gamma$ (chosen to be some representative). By [12, Lem. 3.6] (cf. also [10]), since $\gamma$ is assumed to be nowhere constant, it is a future-directed causal curve. Given $t\in[0,1]$, let $U$ be a neighborhood containing $\gamma(t)$ such that each causal geodesic entirely contained in $U$ is maximizing. W.l.o.g. suppose $t\in(0,1)$ and suppose $\delta>0$ so that $\gamma([t-\delta,t+\delta])\subseteq U$. By uniform convergence, $\gamma_{n}([t-\delta,t+\delta])\subseteq U$ for all large enough $n$. Since $L_{\tau}$ is upper semicontinuous, one may then argue just like in the case of the limit curve theorem (cf. Theorem 2.20) to conclude that $\gamma([t-\delta,t+\delta])$ is maximizing, and in fact $\lim_{n}L_{\tau}(\gamma_{n}|_{[t-\delta,t+\delta]})=L_{\tau}(\gamma|_{[t-\delta,t+\delta]})$ (cf. again the proof of [3, Thm. 2.23])). Since $t$ was arbitrary, we conclude that $\gamma\in\operatorname{Geo}(X)$. Due to the convergence of Lorentzian lengths we just pointed out, we also see that $d_{F}$ and $d_{\Gamma}$ induce the same topology on $\operatorname{Geo}(X)$. ∎
@@
A Lorentzian length space is globally hyperbolic if it is non-totally imprisoning (i.e., $d$-lengths of causal curves in compact sets are uniformly bounded) and $J(x,y)$ is compact for all $x,y\in X$. Globally hyperbolic Lorentzian length spaces are strongly causal, cf. [12, Thm. 3.26].
@@
We may reformulate Theorem 2.27 in the language of equivalence classes as follows: Let $X$ be a globally hyperbolic and locally maximizing Lorentzian length space and $\gamma_{n}\in\operatorname{Geo}(X)$. Suppose that $\gamma_{n}(0)\eqqcolon p_{n}\to p\neq q\leftarrow q_{n}\coloneqq\gamma_{n}(1)$. Then a subsequence of $(\gamma_{n})$ converges in $\operatorname{Geo}(X)$ to $\gamma$. Moreover, if $X$ is regular and $p\ll q$, then $\gamma\in\operatorname{TGeo}(X)$. If each $\gamma_{n}$ is maximizing, then so is $\gamma$.
@@
    $\gamma_{n}^{\prime}(0)\to\gamma^{\prime}(0)$ in $TM$.
@@
Clearly, $(ii)$ implies $(i)$. To see that $(iii)$ implies $(ii)$, observe that if $(p_{n},v_{n})\in TM$ with $(p_{n},v_{n})\to(p,v)$, then the fact that $\gamma_{n}(t)\coloneqq\exp_{p_{n}}(tv_{n})$ converges to $\gamma(t)\coloneqq\exp_{p}(tv)$ in $C^{\infty}_{\mathrm{loc}}$ on the maximal domain of $\gamma$ is a consequence of basic ODE theory, see e.g. [11, Cor. 2.6]. We are left with showing that $(i)$ implies $(iii)$. Suppose $\gamma_{n}\colon[a,b]\to M$ converges uniformly to $\gamma\colon[a,b]\to M$. It is well-known that the exponential map $E\colon\mathcal{D}\subseteq TM\to M^{2}$ is a diffeomorphism from a neighborhood of the zero section in $TM$ onto a neighborhood of the diagonal in $M^{2}$. We may cover $\gamma$ with finitely many convex normal neighborhoods $U_{i}$ such that $\bigcup_{i}U_{i}^{2}$ is contained in the aforementioned neighborhood of the diagonal. For large enough $n$, the $U_{i}$ cover the corresponding portions of $\gamma_{n}$ due to uniform convergence. But then

### v2

@@ 1 Introduction
Our article is structured as follows: In Section 2, we first recall some basics on Lorentzian pre-length spaces used throughout (Subsection 2.1). We then study geodesics and their convergence properties in Subsection $2.$ Due to the lack of arclength parametrization in the generality that we work with, we need to consider equivalence classes of geodesics up to suitable reparametrizations. To this end, in Appendix A, we give a treatment of the Fréchet distance on such equivalence classes of non-stopping curves on a metric spaces that is suitable for our purposes. Our main result in this context is that the set of (equivalence classes of) causal geodesics $\operatorname{Geo}(X)$ is closed in the set of non-stopping curves with respect to the Fréchet distance (see Proposition 2.26).
@@ 2.1 Preliminaries on Lorentzian pre-length spaces
future-directed timelike (resp. causal) curve (past-directed curves are defined analogously) in a Lorentzian pre-length space $X$ is a nowhere constant map $\gamma\colon I=[a,b]\to X$ (i.e., $\gamma$ is not constant on any non-trivial subinterval of II) that is Lipschitz (w.r.t. $d$) and such that for all $t_{1},t_{2}\in I$ with $t_{1}<t_{2}$ we have $\gamma(t_{1})\ll\gamma(t_{2})$ (resp.) $\gamma(t_{1})\leq\gamma(t_{2})$ (causal curves on unbounded intervals can be defined analogously). Its $\tau$-length is defined as
@@
A reparametrization of a future-directed causal curve $\gamma\colon[a,b]\to X$ is a future-directed causal curve $\lambda\colon[c,d]\to X$ such that $\gamma=\lambda\circ\varphi$, where $\varphi\colon[a,b]\to[c,d]$ is surjective and strictly monotonically increasing (hence also continuous). A future-directed causal curve $\gamma\colon[a,b]\to X$ is called a maximizer if $L_{\tau}(\gamma)=\tau(\gamma(a),\gamma(b))$.
In particular, the reparametrized curve $\lambda\circ\varphi$ above is required to itself be Lipschitz. The $\tau$-length of a causal curve is invariant under reparametrizations ([13, Lem. 2.18]). We note that causal curves were defined in [13] as non-constant maps rather than nowhere constant. We choose the current convention for two main reasons, one being that the restriction of a causal curve to any non-trivial subinterval of its domain should again be causal, and the other that causal curves in the standard smooth spacetime setting are indeed nowhere constant. The Fréchet distance as detailed in Appendix A then applies to causal curves.
@@
localizable Lorentzian pre-length space is a Lorentzian pre-length space equipped with well-behaved local time separation functions, see [13, Def. 3.16] for the precise definition. It is called regularly localizable if the maximizers w.r.t. the local time separations (which exist by definition of localizability) are required to be timelike if they connect timelike related points. Regularly localizable Lorentzian pre-length spaces are regular, cf. [13, Thm. 3.18].
@@ 2.2 Convergence of geodesics
A future-directed causal curve $\gamma\colon I\to X$ is a geodesic if for each $t\in I$ there exists a neighborhood $[a,b]$ of $t$ (i.e., $a<t<b$, allowing for equality at the endpoints of II) such that $\gamma|_{[a,b]}$ is a maximizer, i.e., $\tau(\gamma(a),\gamma(b))=L_{\tau}(\gamma|_{[a,b]})$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a Lorentzian pre-length space that is localizable or has a timelike curvature bound (from above or below). Then it is finitely $\tau$-measurable.
@@
The first claim follows from the definition of localizability, cf. [13, Def. 3.16] for the latter. On the other hand, if $X$ has a timelike curvature bound, then $\tau$ is continuous and finite-valued on the product with itself of a comparison neighborhood of any point in $X$. Covering a given causal curve segment with finitely many such neighborhoods establishes finite $\tau$-measurability. ∎
@@
An important class of examples of Lorentzian (pre-) length spaces is furnished by smooth manifolds $M$ equipped with a continuous and causally plain metric $g$ (cf. [13, Prop. 5.8]). To compare the notion of geodesics in Lorentzian manifolds with that of Definition 2.8, let us now extend the validity of [13, Prop. 2.32] from smooth Lorentzian metrics to a large class of continuous ones.
@@
To prove this we employ a sequence of smooth Lorentzian metrics $g_{k}$ with wider lightcones than $g$, approximating $g$ as constructed in [15, Lem. A.1]. Then for the corresponding time separation functions $\tau_{k}$ of $g_{k}$ we have that $\tau$ is the locally uniform and non-increasing limit of the $\tau_{k}$ (see [15, Lem. A.2]). In particular, $L_{\tau_{k}}(\gamma)\geq L_{\tau}(\gamma)$ for all $k$. Let $\varepsilon>0$ and choose a partition $P=\{a=t_{0}<\dots<t_{N}=b\}$ of $[a,b]$ such that
@@
It follows from Proposition 2.13 that (in strongly causal spaces) geodesics in the sense of Definition 2.8 will coincide with geodesics in the sense of solutions of the geodesic equation as long as the regularity of the metric is sufficient to allow for a satisfactory solution theory of these ODEs and, additionally, as long as such solutions are indeed local maximizers of the $g$-length-functional. This certainly is the case for $g\in C^{1,1}$ (cf. [16] or [14]). However, the Hartmann–Wintner-type spacetime $M\coloneqq\mathbb{R}\times(-1,1)\times\mathbb{R}^{n}$ with metric $g\coloneqq-dt^{2}+dx^{2}+(1-|x|^{\lambda})g_{n}$, where $\lambda\in(1,2)$ and $g_{n}$ is the Euclidean metric on $\mathbb{R}^{n}$, is an example of a $C^{1,\lambda-1}$-Lorentzian metric where solutions of the geodesic equations fail to even locally maximize the $g$-length, cf. [20, Ex. 2.4]. In any case, whenever we compare synthetic notions to classical notions in spacetimes we will assume strong causality in order to have a compatible notion of local maximizers.
@@
Note that, unlike [11, Definitions 2.18 & 2.19], we do not require that $(X,d,\ll,\leq,\tau)$ be causally path-connected in our definition of $\leq_{U}$ and, hence, of weakly causally closed. However, we have the following slightly sharpened version of [11, Prop. 2.21]:
@@
    $(X,d,\ll,\leq,\tau)$ is locally causally closed; [^4] Recall from [13, Definition 3.4] that if $(X,d,\ll,\leq,\tau)$ is a Lorentzian pre-length space, then a subset $U\subseteq X$ is causally closed if, for all sequences $(p_{n})_{n}$, $(q_{n})_{n}$ in $U$ such that $p_{n}\leq q_{n}$ for all $n$ with $p_{n}\to p\in\bar{U}$, $q_{n}\to q\in\bar{U}$, we have $p\leq q$. A Lorentzian pre-length space is locally causally closed if every point has a causally closed neighborhood.
@@
Finally, as pointed out in [11, pp. 8], the limit curve theorem [13, Thm. 3.7] holds with the assumption of local weak causal closedness.
@@
From the proof of [13, Thm. 3.7], we deduce that there exists a subsequence $\left(\gamma_{n_{k}}\right)_{k}$ of curves that converge uniformly to a Lipschitz continuous limit curve $\gamma$. Therefore, without loss of generality, we may assume that the sequence of curves $\left(\gamma_{n}\right)_{n}$ converges uniformly to $\gamma$. Since $X$ is locally weakly causally closed, for every $t\in[a,b]$ there exists an open, weakly causally closed neighborhood $U_{t}$ of $\gamma(t)$. Let $a\leq t_{1}<t_{2}\leq b$ be such that $\gamma(t_{1}),\gamma(t_{2})\in U_{t}$. Then, by uniform convergence, there exists an $n_{0}\in\mathbb{N}$ such that, for all $n\geq n_{0}$, we have $\gamma_{n}\!\left(\left[t_{1},t_{2}\right]\right)\subseteq U_{t}$. Since $\gamma_{n}$ are causal, this implies that $\gamma_{n}(t_{1})\leq_{U_{t}}\gamma_{n}(t_{2})$ for all $n\geq n_{0}$. The fact that $U_{t}$ is weakly causally closed then implies that $\gamma(t_{1})\leq_{U_{t}}\gamma(t_{2})$ and, therefore, $\gamma(t_{1})\leq\gamma(t_{2})$. The rest of the proof of part (i) proceeds as in the proof of [13, Lemma 3.6].
@@
Let $(M,g)$ be a strongly causal $C^{1,1}$-spacetime, then it is locally maximizing: By Remark 2.11, the two available notions of timelike (null, causal) geodesics agree. Moreover, by [16, Thm. 6] local maximizers coincide with classical solutions of the geodesic equation. Now let $p\in M$ and $U$ a convex neighborhood containing $p$ (cf. [14, 16]). Then for any two causally related points $q_{1}\leq q_{2}$, $q_{1},q_{2}\in U$, there is a unique causal geodesic connecting them which is entirely contained in $U$. Hence any causal geodesic in $U$ is already maximizing.
@@
The Hartman-Wintner-type spacetimes from [20, Ex. 2.4], which are $C^{1,\alpha}$ for $0<\alpha<1$, are not locally maximizing.
@@
The Lorentzian constant curvature $K$ spaces, which are (the universal coverings of) de-Sitter space, Minkowski space, and anti-de Sitter space of the corresponding radius (cf. Definition 2.3), are locally maximizing: Indeed in the case of $K\geq 0$ this is trivial since causal geodesics are maximizing throughout, and in the case $K<0$ one can consider convex neighborhoods of a given point $x$ which avoid the timelike cut locus $\{\tau(x,.)=\frac{\pi}{\sqrt{-K}}\}$.
@@
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space. We denote by $\operatorname{Geo}(X)$ the set of equivalence classes of future-directed causal geodesic segments $\gamma\colon[0,1]\to X$, where two such segments are called equivalent if and only if there exists a strictly increasing and surjective reparametrization $\varphi\colon[0,1]\to[0,1]$ between them. [^6] See Definition $2.$ We denote by $\operatorname{TGeo}(X)\subseteq\operatorname{Geo}(X)$ the subset of equivalence classes of future-directed timelike geodesic segments. For convenience, we write $\gamma\in\operatorname{Geo}(X)$ for both the equivalence class and some representative.
We consider the following metrics on $\operatorname{Geo}(X)$: $d_{F}$ is the restriction of the Fréchet distance (see Appendix A), and
@@
Let $(X,d,\ll,\leq,\tau)$ be a localizable, locally weakly causally closed, strongly causal, and locally maximizing Lorentzian pre-length space. Then $\operatorname{Geo}(X)$ is a closed subset of the set of nowhere constant curves from $[0,1]$ to $X$ w.r.t. the Fréchet distance $d_{F}$. Moreover, the metric topologies of $d_{F}$ and $d_{\Gamma}$ coincide on $\operatorname{Geo}(X)$.
@@
Let $(\gamma_{n})\subseteq\operatorname{Geo}(X)$ converge to the equivalence class of a nowhere constant path $\gamma\colon[0,1]\to X$. By definition of $d_{F}$, there exists a sequence of representatives (again called $\gamma_{n}$) that converge uniformly to $\gamma$ (chosen to be some representative). By [13, Lem. 3.6] (cf. also [11]), since $\gamma$ is assumed to be nowhere constant, it is a future-directed causal curve. Given $t\in[0,1]$, let $U$ be a neighborhood containing $\gamma(t)$ such that each causal geodesic entirely contained in $U$ is maximizing. W.l.o.g. suppose $t\in(0,1)$ and suppose $\delta>0$ so that $\gamma([t-\delta,t+\delta])\subseteq U$. By uniform convergence, $\gamma_{n}([t-\delta,t+\delta])\subseteq U$ for all large enough $n$. Since $L_{\tau}$ is upper semicontinuous, one may then argue just like in the case of the limit curve theorem (cf. Theorem 2.20) to conclude that $\gamma([t-\delta,t+\delta])$ is maximizing, and in fact $\lim_{n}L_{\tau}(\gamma_{n}|_{[t-\delta,t+\delta]})=L_{\tau}(\gamma|_{[t-\delta,t+\delta]})$ (cf. again the proof of [3, Thm. 2.23]). Since $t$ was arbitrary, we conclude that $\gamma\in\operatorname{Geo}(X)$. Due to the convergence of Lorentzian lengths we just pointed out, we also see that $d_{F}$ and $d_{\Gamma}$ induce the same topology on $\operatorname{Geo}(X)$. ∎
@@
A Lorentzian length space is globally hyperbolic if it is non-totally imprisoning (i.e., $d$-lengths of causal curves in compact sets are uniformly bounded) and $J(x,y)$ is compact for all $x,y\in X$. Globally hyperbolic Lorentzian length spaces are strongly causal, cf. [13, Thm. 3.26].
@@
We may reformulate Theorem 2.27 in the language of equivalence classes as follows: Let $X$ be a globally hyperbolic and locally maximizing Lorentzian length space and $\gamma_{n}\in\operatorname{Geo}(X)$. Suppose that $\gamma_{n}(0)\eqqcolon p_{n}\to p\neq q\leftarrow q_{n}\coloneqq\gamma_{n}(1)$. Then a subsequence of $(\gamma_{n})$ converges in $\operatorname{Geo}(X)$ to some $\gamma$. Moreover, if $X$ is regular and $p\ll q$, then $\gamma\in\operatorname{TGeo}(X)$. If each $\gamma_{n}$ is maximizing, then so is $\gamma$.
@@
    $\gamma_{n}^{\prime}(0)\to\gamma^{\prime}(0)$ in TMTM.
@@
Clearly, $(ii)$ implies $(i)$. To see that $(iii)$ implies $(ii)$, observe that if $(p_{n},v_{n})\in TM$ with $(p_{n},v_{n})\to(p,v)$, then the fact that $\gamma_{n}(t)\coloneqq\exp_{p_{n}}(tv_{n})$ converges to $\gamma(t)\coloneqq\exp_{p}(tv)$ in $C^{\infty}_{\mathrm{loc}}$ on the maximal domain of $\gamma$ is a consequence of basic ODE theory, see e.g. [12, Cor. 2.6]. We are left with showing that $(i)$ implies $(iii)$. Suppose $\gamma_{n}\colon[a,b]\to M$ converges uniformly to $\gamma\colon[a,b]\to M$. It is well-known that the exponential map $E\colon\mathcal{D}\subseteq TM\to M^{2}$ is a diffeomorphism from a neighborhood of the zero section in TMTM onto a neighborhood of the diagonal in $M^{2}$. We may cover $\gamma$ with finitely many convex normal neighborhoods $U_{i}$ such that $\bigcup_{i}U_{i}^{2}$ is contained in the aforementioned neighborhood of the diagonal. For large enough $n$, the $U_{i}$ cover the corresponding portions of $\gamma_{n}$ due to uniform convergence. But then